What is the molarity ℳ (mol/L) of a solution of K2SO4 prepared by dissoving 7.00 g of this compound in 200.0 mL of water?
1
Expert's answer
2018-09-17T09:00:53-0400
Solution.
To find the molarity of the solution it is necessary to use the following formula:
C = m/M*V
m = 7 g
V = 0.2 L
M = 174 g/mol
C = 7/174*0.2 = 0.2 mol/L
Answer:
C = 0.2 mol/L